# LightViz: Autonomous Light-field Surveying and Mapping for Distributed Light Pollution Monitoring

### v1

@@ Abstract
Existing technologies for distributed light-field mapping and light pollution monitoring (LPM) rely on either remote satellite imagery or manual light surveying with single-point sensors such as SQMs (sky quality meters). These modalities offer low-resolution data that are not informative for dense light-field mapping, pollutant factor identification, or sustainable policy implementation. In this work, we propose LightViz – an interactive software interface to survey, simulate, and visualize light pollution maps in real-time. As opposed to manual error-prone methods, LightViz (i) automates the light-field data collection and mapping processes; (ii) provides a platform to simulate various light sources and intensity attenuation models; and (iii) facilitates effective policy identification for conservation.
To validate the end-to-end computational pipeline, we design a distributed light-field sensor suit, collect data on Florida coasts, and visualize the distributed light-field maps. In particular, we perform a case study at St. Johns County in Florida, which has a two-decade conservation program for lighting ordinances. The experimental results demonstrate that LightViz can offer high-resolution light-field mapping and provide interactive features to simulate and formulate community policies for light pollution mitigation. We also propose a mathematical formulation for ‘light footprint’ evaluation, which we integrated into LightViz for targeted LPM in vulnerable communities.
@@ I Introduction
Figure 2: An synopsis of existing sensors and data modalities for LPM is shown [14, 15]: (a) Pros and cons of various sensing modalities; (b) Different twilight types; and (c) luminance meter and SQM for single-point radiance sensing.
@@
For performance validation, we first design and develop a novel light-field sensing module that facilitates long-term radiance sensing (standalone mode) and enables dense light-field sensing in a given area (mobile mode). Compared to traditional single-point measurements with manual surveys, these maps are more accurate, dense, and reliable – enabling long-term LPM [21]. We deploy our sensor suit on a beachfront community in Florida and demonstrate the challenges involved in generating high-resolution distributed light-field maps at scale. These field trials validate the need for LightViz as a user interface to simulate and visualize light-fields for LPM.
Subsequently, we demonstrate the effectiveness of LightViz by conducting a case study in St. Johns County, Florida, which has a two-decade long conservation program [22] for mitigating light pollution along the coast. We configure over $6,000$ streetlights with individual attenuation properties and distribute those across various road types across the county. We find that the light-field map generated by LightViz highlights vulnerable areas, identifies pollutant light sources, and provides an efficient way to track their light footprints over time. Finally, we present various mathematical frameworks on LightViz for community policy identification and prospective solutions for light pollution mitigation.
@@ II-1 Light pollution sensing and mapping solutions
Besides, Unmanned Ground Vehicles (UGVs) [17], on-water Autonomous Surface Vehicles (ASVs) [18], Unmanned Aerial Vehicles (UAVs) [19, 28], and air balloons [29, 30] are used to analyze light pollution in a localized area. UAVs are also deployed to measure light emitted from a particular urban facility such as an outdoor sports complex [31] or a park [32]. In [32], drones are used to extract GPS location and a set of interesting points; human operators then measure light intensity with handheld equipment at those locations. This dynamic assessment process involves carrying heavy equipment such as DSLR cameras [32, 33], illuminance meters [34, 35], and spectrometers [14, 36] – which is labor-intensive and time-consuming. For more efficient and long-term measurement, on-ground cameras and irradiance sensors are deployed at multiple locations. DSLR cameras with fisheye lenses are most commonly used to collect hemispheric photographs [17, 37], which are then processed by Sky Quality Meter (SQM) software for extracting brightness and intensity metrics pertaining to light pollution [16]; see Fig. 2. Both SQM and SQM Lens Ethernet (SQM-LE) technologies make use of the hemispheric observation data to measure night-time brightness in magnitudes per unit area (mag/arc-sec2[14, 38]. These units are qualitatively represented for visualizing spectral power distribution and wavelength sensitivity in a given region.
@@ V-B Vulnerable Community Identification
V-C Community Policy Identification
@@
V-C1 Light Source Placement Strategy
@@
V-C2 Optimal Light Source Type Identification
@@
V-C3 Light Footprint Measurement & Tracking
@@
TABLE IV: Light footprint of the source on four lighting configurations in the vulnerable area 9a.
@@ Acknowledgements
This work is supported in part by the UF research grant $132763$. The authors would like to acknowledge the help from Dr. Blair Witherington of Inwater Research Group Inc. and Rachel Tighe at the Archie Carr Center for Sea Turtle Research. We also thank Dr. Cathi Campbell and Dr. Cynthia Lagueux at UF Department of Biology for their help and support in showcasing our early ideas and prototypes at the 2022-23 Light Pollution Management Workshop events.

### v2

@@ Abstract
Existing technologies for distributed light-field mapping and light pollution monitoring (LPM) rely on either remote satellite imagery or manual light surveying with single-point sensors such as SQMs (sky quality meters). These modalities offer low-resolution data that are not informative for dense light-field mapping, pollutant factor identification, or sustainable policy implementation. In this work, we propose LightViz – an interactive software interface to survey, simulate, and visualize light pollution maps in real-time. As opposed to manual error-prone methods, LightViz (i) automates the light-field data collection and mapping processes; (ii) provides a platform to simulate various light sources and intensity attenuation models; and (iii) facilitates effective policy identification for conservation. To validate the end-to-end computational pipeline, we design a distributed light-field sensor suite, collect data on Florida coasts, and visualize the distributed light-field maps. In particular, we perform a case study at St. Johns County in Florida, which has a two-decade conservation program for lighting ordinances. The experimental results demonstrate that LightViz can offer high-resolution light-field mapping and provide interactive features to simulate and formulate community policies for light pollution mitigation. We also propose a mathematical formulation for ‘light footprint’ evaluation, which we integrated into LightViz for targeted LPM in vulnerable communities. A test-case of the LightViz software release is available at: https://github.com/uf-robopi/LightViz.
@@ I Introduction
In recent years, UAVs equipped with simple RGB cameras have gained prevalence for monitoring urban lighting installations and rendering detailed maps [19, 20]. Targeted aerial data and geospatial maps can help identify light pollution hotspots and quantify upward light emissions accurately. In particular, robust frameworks have been developed for night-ground brightness (NGB) monitoring [18] by combining RGB image indices with ground-based SQM measurements. However, UAV flight times are limited, and night-time image data registration with single-point on-ground sensors requires meticulous calibration [21]. The considerable logistics and time needed for high-resolution measurements further constrain comprehensive LPM at scale [22]. Thus, the image data remains sparse for large-scale deployments; that is, although the generated map is high-resolution, the coverage area per pixel is still high. Such maps are not interactive to investigate mitigation policies in a given community, and are not repeatable for simulating lighting strategies for LPM research.
@@
For performance validation, we first design and develop a novel light-field sensing module that facilitates long-term radiance sensing (standalone mode) and enables dense light-field sensing in a given area (mobile mode). Compared to traditional single-point measurements with manual surveys, these maps are more accurate, dense, and reliable – enabling long-term LPM [24]. We deploy our sensor suit on a beachfront community in Florida and demonstrate the challenges involved in generating high-resolution distributed light-field maps at scale. These field trials validate the need for LightViz as a user interface to simulate and visualize light-fields for LPM. It is important to make the distinction that light-field maps in the computer graphics literature is different [25] – which focuses on the physics of light transport – such as scattering and absorption characteristics within a scene to project ray tracing densities when light propagates through a medium. In the context of light pollution, light-field maps visualize the spatial distribution of artificial light emissions on a geographical scale.
We demonstrate the effectiveness of LightViz by conducting a case study in St. Johns County, Florida, which has a two-decade-long conservation program [26] for mitigating light pollution along the coast. We configure over $6,000$ streetlights with individual attenuation properties and distribute those across various road types across the county. We find that the light-field map generated by LightViz highlights vulnerable areas, identifies pollutant light sources, and provides an efficient way to track their light footprints over time. Finally, we present various mathematical frameworks on LightViz for community policy identification and prospective solutions for light pollution mitigation.
Proposed hypothesis and contributions. The primary hypothesis of this paper is that a lightweight and interactive light-field sensing and mapping system like LightViz can significantly enhance the accuracy, resolution, and scalability of LPM compared to traditional methods. By automating data collection, integrating advanced simulation and attenuation models, and providing real-time high-resolution visualizations in LightViz – we can effectively identify pollution sources, assess their impacts on vulnerable communities, and facilitate the development of informed and sustainable light pollution mitigation policies.
As shown in Fig. 1, LightViz contributes to the LPM technology advancement by: (i) offering end-to-end light-field rendering; (ii) simulating light sources to produce dense maps at significantly higher resolutions; (iii) capturing fine-grained variations in local light-field data; and (iv) addressing latency issues in global map generation. In addition, it facilitates exploring effective light pollution mitigation strategies. users can load existing data and simulate new data to investigate key questions such as: (i) How do the intensity and spatial distributions of light pollution vary across urban, suburban, and rural environments?; (ii) What are the primary factors contributing to light pollution in a given vulnerable area? (iii) How can systematic and effective mitigation strategies be developed to address these issues? Lastly, we introduce a quantitative model to assess the light footprint of individual and collective light sources on a given area. We demonstrate that integrating such a quantitative framework into LightViz can help support precise LPM and targeted mitigation strategies.
@@ II-1 Light pollution sensing and mapping solutions
Figure 2: A synopsis of existing sensors and data modalities for LPM is shown [15, 32]: (a) Pros and cons of various sensing modalities; (b) Different twilight types; and (c) luminance meter and SQM for single-point radiance sensing.
Besides, Unmanned Ground Vehicles (UGVs) [16], on-water Autonomous Surface Vehicles (ASVs) [17], Unmanned Aerial Vehicles (UAVs) [18, 33], and air balloons [34, 35] are used to analyze light pollution in a localized area. UAVs are also deployed to measure light emitted from a particular urban facility such as an outdoor sports complex [36] or a park [37]. In [37], drones are used to extract GPS location and a set of interesting points; human operators then measure light intensity with handheld equipment at those locations. This dynamic assessment process involves carrying heavy equipment such as DSLR cameras [37, 38], illuminance meters [39, 40], and spectrometers [15, 41] – which is labor-intensive and time-consuming. For more efficient and long-term measurement, on-ground cameras and irradiance sensors are deployed at multiple locations. DSLR cameras with fisheye lenses are most commonly used to collect hemispheric photographs [16, 42], which are then processed by Sky Quality Meter (SQM) software for extracting brightness and intensity metrics pertaining to light pollution [14]. Fig. 2a provides an index of different modalities highlighting that digital cameras dominate in the spectral range, while illuminance meters are the most user-friendly devices. For nighttime, different twilight types are distinguished based on time, as shown in Fig. 2b. This is a key factor for collecting light pollution data, as the natural background light varies at different times. Both SQM and SQM Lens Ethernet (SQM-LE) technologies make use of the hemispheric observation data to measure night-time brightness in magnitudes per unit area (mag/arc-sec2[15, 43]. These units are qualitatively represented for visualizing spectral power distribution and wavelength sensitivity in a given region.
@@ IV-D Light-Field Map Rendering
An operational outline of the LightViz software is shown in Fig. 7. Once launched, users can load pre-compiled data, add new light sources and their configurations, select interpolation algorithms, and generate light-field maps in real-time. More importantly, users can interact with the map by zooming in and simulating various light-field layouts for policy identification in the selected areas.
Figure 7: An outline of the user commands and interaction flow for the proposed LightViz software interface is shown.
IV-E Attenuation configuration and measurement guidelines
Estimating accurate light attenuation curves is challenging due to the dependencies on complex physical properties of light-particle interaction in a given environment [63]. We follow the general practice in the literature [62] to simplify this process using curve fitting. As shown earlier in Fig. 6, we adopt a quadratic attenuation pattern with two parameters estimated by measuring light intensities at various distances from the source using an SQM. Note that the integration of any other physically accurate model to LightViz is trivial; users can enter exact values (if known) or simulate various configurations of light source types, e.g., narrow emission spectra (LED lights), broader spectra (high-pressure sodium lights), broader spectra with rapid attenuation (incandescent bulbs), etc.
Besides, the number and arrangement of measurement points are also critical for accurate interpolation. The resolution of the interpolation depends on the number of points, which users can adjust based on their requirements. To ensure meaningful results, at least three non-collinear points are minimally required, although $20$ or more points are ideal to generate meaningful results. LightViz includes a built-in validation to check for sufficient and properly arranged points, providing user feedback when necessary (see Fig. 7). This approach balances practical usability with flexibility, enabling users to model light attenuation effectively under different scenarios.
@@ V-A Light Source Layouts and Types
In our field experimental deployments, we demonstrated how our low-power compact sensor nodes can be deployed for distributed light-field estimation and mapping. Our LightViz simulator allows us to visualize the deployed sensor data, and validate the accuracy of various geospatial interpolation algorithms. However, it is logistically challenging to conduct large-scale deployments across cities or counties for comprehensive validations. Once deployed, it also becomes challenging to alter light source types and intensity profiles to simulate various scenarios.
@@ V-B Vulnerable Community Identification
V-C Comparison with UAV-based aerial maps
UAVs can effectively capture targeted aerial imagery to identify light pollution hotspot maps from night-ground brightness (NGB) data [19, 20]. However, their utility is limited in county-scale map generation for long-term LPM due to their short flight times, geospatial image calibration with on-ground sensors, and other logistical requirements [22, 21]. Thus, community pollution tracking and effective policy identification become challenging at scale [21, 18].
In contrast, LightViz offers light source (type and placement) simulation and interactive features to generate high-resolution maps across varying scales. These capabilities facilitate the development of light pollution mitigation policies in a given community. To illustrate the differences, we compare LightViz with drone-gonio-photometer (DGPM) [20], an integrated technology that utilizes drones to plan trajectories and collect light pollution data. As shown in Fig. 10, LightViz offers broader coverage compared to DGPM and provides seamless transitions between large-scale and small-scale views. It also shows that although the DGPM map is high-resolution, the coverage area per pixel is still high. Moreover, the aerial maps are static and are not repeatable for simulating various lighting strategies and planning, as well as policymaking in large urban areas.
Figure 10: (a) Photo mosaic of an interchange from [20]; (b) the DGPM flight path over the interchange; (c) LightViz simulation of light pollution in St. Johns County; and (d) a zoom-in version illustrating seamless transitions between large-scale and small-scale views.
V-D Community Policy Identification
@@
V-D1 Light Source Placement Strategy
@@
V-D2 Optimal Light Source Type Identification
@@
V-D3 Light Footprint Measurement & Tracking
@@
TABLE IV: Light footprint of the source on four lighting configurations in the vulnerable area 11a.
@@
V-E Future works
LightViz offers several key features that make it a powerful tool for LPM, enabling high-resolution light-field mapping and real-time data collection with geospatial visualization. It also supports customizable light attenuation models, allowing users to simulate various lighting conditions and analyze their impacts. We also highlighted a particularly valuable feature for future researchers, which is to measure and track light footprints systematically. These capabilities make LightViz an essential tool for advancing sustainable lighting solutions and ecological conservation efforts. To this end, we highlight the following research directions that will benefit the advances in LPM technology of the next generation.
  1. $1$)
    AI-driven models for light-field estimation and interpolation: By leveraging AI and machine learning algorithms trained on simulated data generated by LightViz, robust predictive models can be formulated to identify pollution trends and emerging hotspots before they become severe. Deep learning-based pipelines can also be developed to automate the classification of pollutant sources and assess their impact on nearby communities. Furthermore, real-time anomaly detection algorithms can be implemented to monitor sudden changes in lighting conditions, helping urban planners and environmentalists respond proactively.
  2. $2$)
    Refined mathematical modeling of light footprint assessment: Further research on LightViz will facilitate the development of a more comprehensive mathematical framework for light footprint measurement and tracking. This could involve the integration of parametric and non-parametric modeling techniques to enhance the accuracy of footprint estimations. Incorporating community-specific pollutant parameters could also improve the precision of light pollution impact predictions for more sustainable urban planning. We envision that our formulations present in this paper will serve as a good baseline, with which future research works can converge to more improved formulations of light footprint assessment and tracking.
  3. $3$)
    Policy simulation and decision support systems: Another significant extension of LightViz will be the development of an automated policy simulation and decision support system. Such a system will allow city planners and policymakers to simulate various lighting policies, such as streetlight dimming schedules, fixture replacements, and zoning regulations, to evaluate their impact before implementation. Additionally, an AI-powered recommendation engine can be developed to suggest optimized lighting configurations tailored to specific urban environments.
These advancements will enable LightViz to transition from a human-operated mapping tool to a dynamic forecasting system that facilitates intelligent decision-making for sustainable lighting practices. Such an automated tool, in the form of a citizen app, will help engage local communities to improve awareness and collect extensive data for more effective light pollution mitigation.
@@ Acknowledgements
This work is supported in part by the UF research grant #$132763$. The authors would like to acknowledge the help from Dr. Blair Witherington of Inwater Research Group Inc. and Rachel Tighe at the Archie Carr Center for Sea Turtle Research. We also thank Dr. Cathi Campbell and Dr. Cynthia Lagueux at UF Department of Biology for their help and support in showcasing our early ideas and prototypes at the 2022-23 Light Pollution Management Workshop events.